OpenStudy (anonymous):

Write an expression in simplest form for the volume of a cube with a side of 11x⁶ units.

OpenStudy (anonymous):

You're finding the Volume of a cube, the formula is \[V=Sides ^3\]

OpenStudy (anonymous):

So, \[V=(11x^6)^3\]

OpenStudy (anonymous):

Using rule \[(mn)^p=m^pn^p\] You get \[V=11^3(x^6)^3\] Simplify

OpenStudy (jdoe0001):

 \(\bf \textit{volume of the cube} = 11x^6 \times 11x^6 \times 11x^6 \implies (11x^6)^3\)